Solve using Polya’s four-step problem solving strategies.

2. According to a student survey on keyboard light preference, 44% liked white, 26% liked blue and 30% liked green. 8% liked both white and green, 12% liked both white and blue, 10% liked both green and blue. If 84% of them liked at least one keyboard light color, what percentage of people liked all three? 

 

 


1
Expert's answer
2021-02-25T23:31:25-0500

If 44% liked white, 26% liked blue and 30% liked green, and 84% liked at least one keyboard light color then (44 + 26 + 30) - 84 = 16% liked more than 1 color.

As 8% liked both white and green, 12% liked both white and blue, 10% liked both green and blue, so the total number is 8 + 12 + 10 = 30%, but only 16% liked more than 1 color.

So, 30 - 16 = 14% liked all three colors.